# Training-Conditional Coverage Bounds under Covariate Shift

### v3

@@ Abstract
Conformal prediction methodology has recently been extended to the covariate shift setting, where the distribution of covariates differs between training and test data. While existing results ensure that the prediction sets from these methods achieve marginal coverage above a nominal level, their coverage rate conditional on the training dataset—referred to as training-conditional coverage—remains unexplored. In this paper, we address this gap by deriving upper bounds on the tail of the training-conditional coverage distribution, offering probably approximately correct (PAC) guarantees for these methods. Our results quantify the relationship between the quality of the prediction sets and the severity of distributional changes, and can potentially be used to compute more efficient prediction sets.
@@ 1 Introduction
Conformal prediction is a framework for constructing distribution-free and model-agnostic predictive confidence regions under the exchangeability assumption for the training and test samples Vovk et al. (2005) (also see Shafer and Vovk (2008); Vovk et al. (2009); Vovk (2012)). Let $((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(X_{\text{test}},Y_{\text{test}}))$ denote a tuple of exchangeable data points, consisting of a training sequence of $n$ samples ${\bf D}_{n}:=((X_{1},Y_{1}),\ldots,(X_{n},Y_{n}))\in(\mathcal{X}\times\mathcal{Y})^{n}$ and one test sample $(X_{\text{test}},Y_{\text{test}})\in\mathcal{X}\times\mathcal{Y}$. For a fixed error rate level $\alpha$, the conformal prediction framework, provides a prediction set $\hat{C}_{\alpha}(X_{\text{test}})$ for $Y_{\text{test}}$ that satisfies
$\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{C}_{\alpha}(X_{\text{test}}))\geq 1-\alpha,~{}$ (1)
@@
$\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{\mathcal{C}}_{\alpha}(X_{\text{test}})|X_{\text{test}})\geq 1-\alpha.$
@@
$P_{e}({\bf D}_{n}):=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}(X_{\text{test}})|{\bf D}_{n}).$
@@
$\operatorname{\mathbb{P}}(P_{e}({\bf D}_{n})\geq\alpha)\leq\delta$
@@
is established in Bian and Barber (2023). They have also shown that distribution-free training-conditional guarantees for full conformal and jackknife+ methods are impossible without further assumptions; in particular, they conjectured that a certain form of algorithmic stability is needed for full conformal and jackknife+. Recently, Liang and Barber (2023) proposed conditional coverage bounds for jackknife+ and full conformal prediction sets under the assumption that the training algorithm is symmetric and satisfies certain stability conditions (see Section 4 for more details). In this line of research, samples are assumed to be i.i.d., which is not only exchangeable but also ergodic and admits some nice concentration properties. However, this assumption can be violated in the application. In particular, the input data distribution during deployment can differ from the distribution observed during training. This phenomenon is called distribution shift and it is a crucial problem in trustworthy machine learning (see Section 2.3). In this regard, split and full conformal prediction methods as the central methods for distribution-free uncertainty quantification of black-box models have recently been extended to handle a popular type of distribution shift called covariate shift Tibshirani et al. (2019). In the covariate shift setting, the distribution of covariates in the test data differs from the one observed in the training data, but, the conditional distribution of the response given the features remains the same across the training and test populations. A similar extension has been made for the jackknife+ method Prinster et al. (2022). However, these results primarily focus on marginal coverage. Although the weighted conformal prediction is guaranteed to keep the marginal coverage rate above the nominal level, it often reduces the concentration of the training-conditional coverage rate $\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{C}_{\alpha}(X_{\text{test}})|{\bf D}_{n})$ (i.e., coverage rate conditional on the training data ${\bf D}_{n}$) around the nominal level $1-\alpha$. In particular, the weighting scheme leads to heavier tails for the training-conditional coverage rate. This phenomenon is illustrated in Figure 1, where ordinary and weighted conformal methods with $80\%$ target coverage rate ($\alpha=0.2$) are used for constructing prediction intervals under exchangeability and distribution shift settings, respectively. The distribution shift is introduced artificially via resampling according to an exponential tilt as in Section 2.3 from Tibshirani et al. (2019). It can be observed that the tails of the training conditional coverage get heavier when the weighting scheme is used to handle the distribution shift. In this paper, we explore the quality of the weighted conformal prediction regions by computing upper bounds on the tails of the training-conditional coverage distribution, quantifying the relationship between the tail behavior and the extent of distribution change. In other words, we examine the efficiency of weighted conformal prediction methods for a generic training dataset under distribution shift.
Figure 1: Histograms of the training-conditional coverage rates are presented for four datasets from the UCI Machine Learning Repository: Wine Quality (top left), Abalone (top right), Concrete Compressive Strength (bottom left), and Combined Cycle Power Plant (bottom right). Models were trained using 1000 samples and an RKHS regression model with a Gaussian kernel.
@@
We present concentration bounds of the training-conditional errors for weighted jackknife+ (JAW) Prinster et al. (2022), full, and split conformal methods. Regarding the training algorithm, no assumption is made for the split conformal method. However, full conformal and jackknife+ methods are analyzed under the assumption of uniform stability as explained below. Let $\hat{\mu}_{{\bf D}_{n}}$ denote the predictor function estimated using the training data ${\bf D}_{n}$. A training algorithm is called uniformly stable if,
@@
with $\beta=O(1/n)$ for any two datasets $({\bf D}_{n}$, ${\bf D}^{\prime}_{n})$ differing in one (training) data sample Bousquet and Elisseeff (2002). This is a stronger notion of algorithmic stability than the $(m,n)$-stability assumed in Liang and Barber (2023). Nevertheless, uniform stability is satisfied by a class of regression models known as RKHS regression Bousquet and Elisseeff (2002), i.e., regularized empirical risk minimization over an RKHS Paulsen and Raghupathi (2016); Schölkopf and Smola (2002). Examples of reproducing kernel Hilbert spaces are certain Sobolev spaces of smooth functions Wahba (1990).
@@ 2.1 Full conformal and split conformal
Define the shorthand ${\bf D}_{n}\cup(x,y):=((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(x,y))$ and let $\hat{\mu}_{(x,y)}:=T({\bf D}_{n}\cup(x,y))$ denote the regression function obtained by running the training algorithm $T$ on ${\bf D}_{n}\cup(x,y)$. Define the score function $s(x^{\prime},y^{\prime};\mu):=f(\mu(x^{\prime}),y^{\prime})$ via some arbitrary (measurable) cost function $f:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ and predictor function $\mu:\mathcal{X}\to\mathcal{Y}$. For instance, $s(x^{\prime},y^{\prime};\mu)=|y^{\prime}-\mu(x^{\prime})|$ when $f(y,y^{\prime})=|y-y^{\prime}|$. Define the (training) multiset $\mathcal{D}_{n}:=\{(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}:i\in[n]\}$ where $[n]:=\{1,2,...,n\}$, and let
$\mathcal{S}(D;\mu):=\{s(x^{\prime},y^{\prime};\mu):(x^{\prime},y^{\prime})\in D\},\quad D\in\cup_{n\geq 1}(\mathcal{X}\times\mathcal{Y})^{n}.$
Observe that if $T$ is symmetric, i.e., $T(D)$ is invariant under permutations of the elements of the training data $D$, then the elements of $\mathcal{S}(\mathcal{D}_{n}\cup(X_{\text{test}},Y_{\text{test}});\hat{\mu}_{(X_{\text{test}},Y_{\text{test}})})$ are exchangeable. Therefore,
@@
where $\hat{F}^{-1}_{S}(\cdot)$ denotes the empirical quantile function with respect to $S$. Thus,
$\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{C}_{\alpha}^{\text{full}}(X_{\text{test}}))\geq 1-\alpha,$
@@
$\hat{C}^{\text{full}}_{\alpha}(x)$ =$=\{y:s(x,y;\hat{\mu}_{(x,y)})\leq\hat{F}^{-1}_{\mathcal{S}(\mathcal{D}_{n}\cup(x,y);\hat{\mu}_{(x,y)})}(1-\alpha)\}$
@@
It is well-known that this approach can be computationally intensive when $\mathcal{Y}=\mathbb{R}$ since to find out whether $y\in\hat{C}^{\text{full}}_{\alpha}(x)$ one needs to train the model with the dataset including $(x,y)$. One simple way to alleviate this issue is to split the data into training and calibration datasets, namely $\mathcal{D}_{n}=\mathcal{D}^{\text{train}}\cup\mathcal{D}^{\text{cal}}$. Without loss of generality (when splitting), we assume ${\bf D}^{\text{train}}=((X_{1},Y_{1}),\ldots,(X_{n_{0}},Y_{n_{0}}))$ and ${\bf D}^{\text{cal}}=((X_{n_{0}+1},Y_{n_{0}+1}),\ldots,(X_{n},Y_{n}))$. In the split conformal method, one first finds the regression predictor $\hat{\mu}=T({\bf D}^{\text{train}})$ and treats $\hat{\mu}$ as fixed. Note that the elements of $\mathcal{S}(\mathcal{D}^{\text{cal}}\cup(X_{\text{test}},Y_{\text{test}});\hat{\mu})$ are exchangeable. Hence, we get
@@
$\hat{C}^{\text{split}}_{\alpha}(x)$ =$=\left\{y:s(x,y;\hat{\mu})\leq\hat{F}^{-1}_{\mathcal{S}(\mathcal{D}^{\text{cal}};\hat{\mu})\cup\{\infty\}}(1-\alpha)\right\}$ (5)
@@ 2.2 Jackknife+
$\mathcal{S}^{-}(x)$ =$=\{\hat{\mu}^{-i}(x)-|Y_{i}-\hat{\mu}^{-i}(X_{i})|:i\in[n]\}\cup\{-\infty\},$
$\mathcal{S}^{+}(x)$ ={${}.\displaystyle=\{\hat{\mu}^{-i}(x)+|Y_{i}-\hat{\mu}^{-i}(X_{i})|:i\in[n]\}\cup\{\infty\}.$
@@ 2.3 Conformal prediction under distribution shift
[…]
$Z_{\text{test}}$ $:=(X_{\text{test}},Y_{\text{test}})\sim Q_{X}\times P_{Y|X}=:Q\quad\text{ test data.}$.
@@
Similarly, to extend the jackknife+ method, Prinster et al. (2022) proposes replacing $\hat{F}_{\mathcal{S}^{-}}$ and $\hat{F}_{\mathcal{S}^{+}}$ by their corresponding weighted versions where $w_{\text{test}}$ is associated to $\infty$- ∞ and $\infty$, respectively.
@@ 3.1 Split conformal
The split conformal training-conditional coverage guarantee under the i.i.d. setting has been studied in Vovk (2012); Bian and Barber (2023). The proof in this setting relies on the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality. The following theorem concerns training-conditional coverage guarantee under covariate shift, where the extension is made by developing a weighted version of the DKW inequality. Let ${\bf D}_{n}=(Z_{1},\ldots,Z_{n})\in(\mathcal{X}\times\mathcal{Y})^{n}$ denote an $n$-tuple of data points containing both the training and calibration data sets. Define the training conditional probability of error
$P_{e}^{\text{split}}({\bf D}_{n})=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{split}}_{\alpha}(X_{\text{test}})\Big{|}{\bf D}_{n}\right).$ (7)
@@
Let $m<n$ denote the size of the calibration data set. Assume $Q\ll P$ and $dQ/dP\leq B<\infty$. Then,
@@
Relaxing the assumption $dQ/dP\leq B$ to $\|dQ/dP\|_{P,2}\leq B$ via Lemma 2, we get
@@
To assess the robustness of weighted conformal prediction methods, we compute training-conditional error bounds with no assumptions on $dQ/dP$. The proof follows from Lemma 4 and the proof of Theorem 1.
@@
Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
$P_{e}^{\text{split}}({\bf D}_{n})<\alpha+3\delta+4\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>m\right)+C^{\prime}\sqrt{\frac{B}{m^{\prime}}}+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right),$
@@
$-\frac{8}{m\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq m\right\}\right)},$

### v4

@@ Abstract
Conformal prediction methodology has recently been extended to the covariate shift setting, where the distribution of covariates differs between training and test data. While existing results ensure that the prediction sets from these methods achieve marginal coverage above a nominal level, their coverage rate conditional on the training dataset—referred to as training-conditional coverage—remains unexplored. In this paper, we address this gap by deriving upper bounds on the tail of the training-conditional coverage distribution, offering probably approximately correct (PAC) guarantees for these methods. Our results characterize the reliability of the prediction sets in terms of the severity of distributional changes and the size of the training dataset.
@@ 1 Introduction
Conformal prediction is a framework for constructing distribution-free and model-agnostic predictive confidence regions under the exchangeability assumption for the training and test samples (Vovk et al., 2005) (also see Shafer and Vovk (2008); Vovk et al. (2009); Vovk (2012)). Let $((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(X_{\text{test}},Y_{\text{test}}))$ denote a tuple of exchangeable data points, consisting of a training sequence of $n$ samples ${\bf D}_{n}:=((X_{1},Y_{1}),\ldots,(X_{n},Y_{n}))\in(\mathcal{X}\times\mathcal{Y})^{n}$ and one test sample $(X_{\text{test}},Y_{\text{test}})\in\mathcal{X}\times\mathcal{Y}$. For a fixed error rate level $\alpha$, the conformal prediction framework, provides a prediction set $\hat{C}_{\alpha}(X_{\text{test}})$ for $Y_{\text{test}}$ that satisfies
$\operatorname{\mathbb{P}}\left(Y_{\text{test}}\in\hat{C}_{\alpha}(X_{\text{test}})\right)\geq 1-\alpha,$~ (1)
@@
$\operatorname{\mathbb{P}}\left(Y_{\text{test}}\in\hat{\mathcal{C}}_{\alpha}(X_{\text{test}})|X_{\text{test}}\right)\geq 1-\alpha.$
@@
$P_{e}({\bf D}_{n}):=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}(X_{\text{test}})|{\bf D}_{n}\right).$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})\geq\alpha\right)\leq\delta$
@@
is established in Bian and Barber (2023). Additionally, Liang and Barber (2023) proposed training-conditional coverage bounds for jackknife+ and full conformal prediction sets under the assumption that the training algorithm is symmetric and satisfies certain stability conditions (see Section 4 for more details). In this line of research, samples are assumed to be i.i.d., which is not only exchangeable but also ergodic and admits some nice concentration properties. However, this assumption can be violated in the application. In particular, the input data distribution during deployment can differ from the distribution observed during training. This phenomenon is called distribution shift and it is a crucial problem in trustworthy machine learning (see Section 2.3). In this regard, split and full conformal prediction methods as the central methods for distribution-free uncertainty quantification of black-box models have recently been extended to handle a popular type of distribution shift called covariate shift (Tibshirani et al., 2019). In the covariate shift setting, the distribution of covariates in the test data differs from the one observed in the training data, but, the conditional distribution of the response given the features remains the same across the training and test populations. A similar extension has been made for the jackknife+ method Prinster et al. (2022).
Despite these significant progress on handling distribution shift, they primarily focus on marginal coverage. Although the weighted conformal prediction methods (proposed in Tibshirani et al. (2019) and Prinster et al. (2022)) are guaranteed to keep the marginal coverage rate above the nominal level, they often reduce the concentration of the training-conditional coverage rate $\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{C}_{\alpha}(X_{\text{test}})|{\bf D}_{n})$ (i.e., coverage rate conditional on the training data ${\bf D}_{n}$) around the nominal level $1-\alpha$. In particular, the weighting scheme leads to heavier tails for the training-conditional coverage rate. This phenomenon is illustrated in Figure 1, where ordinary and weighted conformal methods with $80\%$ target coverage rate ($\alpha=0.2$) are used for constructing prediction intervals under exchangeability and distribution shift settings, respectively. The distribution shift is introduced artificially via resampling according to an exponential tilt as in Section 2.3 from Tibshirani et al. (2019). It can be observed that the tails of the training conditional coverage get heavier when the weighting scheme is used to handle the distribution shift. In this paper, we explore the quality of the weighted conformal prediction sets by computing upper bounds on the tails of the training-conditional coverage distribution, quantifying the relationship between the tail behavior and the extent of distribution change. In other words, we examine the efficiency of weighted conformal prediction methods for a generic training dataset under distribution shift.
@@
We present concentration bounds of the training-conditional coverage for weighted jackknife+ (JAW) (Prinster et al., 2022), full, and split conformal methods. Regarding the training algorithm, no assumption is made for the split conformal method. However, full conformal and jackknife+ methods are analyzed under the assumption of uniform stability as explained below. Let $\hat{\mu}_{{\bf D}_{n}}$ denote the predictor function estimated using the training data ${\bf D}_{n}$. A training algorithm is called uniformly stable if,
@@
with $\beta=O(1/n)$ for any two datasets $({\bf D}_{n}$, ${\bf D}^{\prime}_{n})$ differing in one (training) data sample (Bousquet and Elisseeff, 2002). This is a stronger notion of algorithmic stability than the $(m,n)$-stability assumed in Liang and Barber (2023). Nevertheless, uniform stability is satisfied by a class of regression models known as RKHS regression (Bousquet and Elisseeff, 2002), i.e., regularized empirical risk minimization over an RKHS (Paulsen and Raghupathi, 2016; Schölkopf and Smola, 2002). Examples of reproducing kernel Hilbert spaces are certain Sobolev spaces of smooth functions (Wahba, 1990).
Figure 1: Histograms of the training-conditional coverage rates are presented for four datasets from the UCI Machine Learning Repository: Wine Quality (top left), Abalone (top right), Concrete Compressive Strength (bottom left), and Combined Cycle Power Plant (bottom right). See Appendix E for details of this simulation study.
@@ 2.1 Full conformal and split conformal
Define the shorthand ${\bf D}_{n}\cup(x,y):=((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(x,y))$ and let $\hat{\mu}_{(x,y)}:=T({\bf D}_{n}\cup(x,y))$ denote the regression function obtained by running the training algorithm $T$ on ${\bf D}_{n}\cup(x,y)$. Define the score function $s(x^{\prime},y^{\prime};\mu):=f(\mu(x^{\prime}),y^{\prime})$ via some arbitrary (measurable) cost function $f:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ and predictor function $\mu:\mathcal{X}\to\mathcal{Y}$. For instance, $s(x^{\prime},y^{\prime};\mu)=|y^{\prime}-\mu(x^{\prime})|$ when $f(y,y^{\prime})=|y-y^{\prime}|$. For the data multiset $\mathcal{D}_{n}:=\{(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}:i\in[n]\}$ where $n]:=\{1,2,...,n\}$, define
$\mathcal{S}(\mathcal{D}_{n};\mu):=\{s(x^{\prime},y^{\prime};\mu):(x^{\prime},y^{\prime})\in\mathcal{D}_{n}\}.$
Observe that if $T$ is symmetric, i.e., $T({\bf D})$ is invariant under permutations of the elements of $\bf D$} for any data tuple $\bf D$}, then the elements of $\mathcal{S}(\mathcal{D}_{n}\cup(X_{\text{test}},Y_{\text{test}});\hat{\mu}_{(X_{\text{test}},Y_{\text{test}})})$ are exchangeable. Therefore,
@@
where $\hat{F}^{-1}_{S}(\cdot)$ denotes the empirical quantile function of $S$. Thus,
$\operatorname{\mathbb{P}}\left(Y_{\text{test}}\in\hat{C}_{\alpha}^{\text{full}}(X_{\text{test}})\right)\geq 1-\alpha,$
@@
$\hat{C}^{\text{full}}_{\alpha}(x)$ $=\{y:s(x,y;\hat{\mu}_{(x,y)})\leq\hat{F}^{-1}_{\mathcal{S}(\mathcal{D}_{n}\cup(x,y);\hat{\mu}_{(x,y)})}(1-\alpha)\}$
@@
[…]
@@
$\hat{C}^{\text{split}}_{\alpha}(x)$ $=\left\{y:s(x,y;\hat{\mu})\leq\hat{F}^{-1}_{\mathcal{S}(\mathcal{D}^{\text{cal}};\hat{\mu})\cup\{\infty\}}(1-\alpha)\right\}$ (5)
@@ 2.2 Jackknife+
$\mathcal{S}^{-}(x)$ $=\{\hat{\mu}^{-i}(x)-|Y_{i}-\hat{\mu}^{-i}(X_{i})|:i\in[n]\}\cup\{-\infty\},$
$\mathcal{S}^{+}(x)$ $=\{\hat{\mu}^{-i}(x)+|Y_{i}-\hat{\mu}^{-i}(X_{i})|:i\in[n]\}\cup\{\infty\}.$
@@ 2.3 Conformal prediction under distribution shift
$Z_{i}$ $:=(X_{i},Y_{i})\overset{\text{ i.i.d. }}{\sim}P_{X}\times P_{Y|X}=:P\quad\text{ training data;}$
$Z_{\text{test}}$ $:=(X_{\text{test}},Y_{\text{test}})\sim Q_{X}\times P_{Y|X}=:Q\quad\text{ test data.}$
@@
Similarly, to extend the jackknife+ method, Prinster et al. (2022) proposes replacing $\hat{F}_{\mathcal{S}^{-}}$ and $\hat{F}_{\mathcal{S}^{+}}$ by their corresponding weighted versions where $w_{\text{test}}$ is associated to -$\infty$ and $\infty$, respectively.
@@ 3.1 Split conformal
The training-conditional coverage guarantee for split conformal method under the i.i.d. setting has been studied in Vovk (2012); Bian and Barber (2023). The proof in this setting relies on the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality. The following theorem concerns training-conditional coverage guarantee under covariate shift, where the extension is made by developing a weighted version of the DKW inequality. Let ${\bf D}_{n}=(Z_{1},\ldots,Z_{n})\in(\mathcal{X}\times\mathcal{Y})^{n}$ denote an $n$-tuple of data points containing both the training and calibration data sets. Define the training conditional probability of error as
$P_{e}^{\text{split}}({\bf D}_{n}):=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{split}}_{\alpha}(X_{\text{test}})\Big|{\bf D}_{n}\right).$ (7)
@@
Let $m<n$ denote the size of the calibration data set. Assume that $Q$ is absolutely continuous with respect to $P$ ($Q\ll P$) and $dQ/dP\leq B<\infty$. Then,
@@
Relaxing the assumption $dQ/dP\leq B$ to $\|dQ/dP\|_{P,2}:=\int(dQ/dP)^{2}~dP\leq B$ via Lemma 2, we get
@@
To assess the robustness of weighted conformal prediction methods, we compute training-conditional error bounds with a very mild assumption on $dQ/dP$. The proof of this theorem follows from Lemma 4 and the proof of Theorem 1.
@@
Assume that $Q$ is absolutely continuous with respect to $P$ ($Q\ll P$) and that ${dQ}/{dP}<\infty$ Q-a.s. Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
$P_{e}^{\text{split}}({\bf D}_{n})<\alpha+3\delta+4\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>m\right)+2C\sqrt{\frac{B}{m^{\prime}}}+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right),$
@@
$-\frac{8}{m\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}(Z_{\text{calib}})\right)^{2}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq m\right\}\right)},$